OpenStudy (anonymous):

abs(7x+5)=abs(7x-9) solve for x

OpenStudy (anonymous):

abs=absolute value?

OpenStudy (anonymous):

yes gandalf the white

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

I prefer Gandalf the Green... until Ievel up and my profile color changes

OpenStudy (anonymous):

\[\left| 7x+5 \right|=\left| 7x-9 \right|\]

OpenStudy (anonymous):

thats it but i need x

OpenStudy (anonymous):

I know... working on it...

OpenStudy (anonymous):

sorry, it's been a while... I thought that you had to change all the negative signs to positive and solve. But that doesn't seem to be working :(

OpenStudy (zarkon):

\[\frac{2}{7}\]

OpenStudy (anonymous):

how u solve, zarkon?

OpenStudy (zarkon):

solve 7x+5=-(7x-9)

OpenStudy (anonymous):

... why? what did u do?

OpenStudy (zarkon):

there are 4 possibilities but only one gives a valid solution

OpenStudy (zarkon):

(7x+5)=(7x-9) (7x+5)=-(7x-9) -(7x+5)=(7x-9) -(7x+5)=-(7x-9)

OpenStudy (zarkon):

I guess 2 give a valid solution

OpenStudy (anonymous):

how do u know that the others are wrong?

OpenStudy (zarkon):

solve them

OpenStudy (zarkon):

7x+5=7x-9 5=-9

OpenStudy (anonymous):

got it.

OpenStudy (anonymous):

one way to do these is to square both sides: 49x^2 +70x +25 = 49x^2 - 126x + 81 196x = 56 x = 56/196

OpenStudy (anonymous):

this checks out LHS = 7, RHS = -7 - absolute value 7

OpenStudy (anonymous):

-same as zarkon as 56/186 = 2/7